# Beyond Reference-Based Metrics: Analyzing Behaviors of Open LLMs on Data-to-Text Generation

### v1

@@ 4.1 GPT-4-based Evaluation
LLM-based metrics can be customized for a particular task without the need for training data. For our experiments, we employ a metric based on GPT-4 (gpt-4-1106-preview, OpenAI, 2023a), which was shown to be superior in following fine-grained instructions compared to other LLMs, and to be having high correlations with human judgement on evaluating generated texts Zhao et al. (2023); Sottana et al. (2023); Kocmi and Federmann (2023a, b).
$\mathcal{E}_{\text{gpt}}$ is instantiated using a prompt and a system message describing the task. We instruct the model to produce a JSON output with sequentially ordered errors using the following format:
@@
Note that we require that the model first generates the reason for the error. Besides that we get a free-form text reasoning for each error without excessive extra costs, our cursory observations suggest that this leads to more precise outputs.
Concerning the alignment of the model outputs with the original text, we perform matching on TEXT_SPAN. We ensure that the model response is a valid JSON using OpenAI’s response_format parameter. See Appendix B for more details about the metric, including the prompt and the system message.
4.2 Human-based Evaluation
An automatic metric based on a closed LLM makes the evaluation potentially non-reproducibile and biased Kocmi and Federmann (2023a); Wang et al. (2023b), for which we compensate by obtaining annotations from human annotators.
For the human annotation metric $\mathcal{E}_{\text{hum}}$, we prepared a custom web interface based on TabGenie Kasner et al. (2023a), where an annotator is able to annotate a text span with a selected error category. We created custom visualizations for each data format. Unlike with $\mathcal{E}_{\text{gpt}}$, we did not ask the crowdworkers for free-form reasoning about the errors since that would make the annotation more complex.
We hired annotators on the Prolific crowdsourcing platform. [^5] https://prolific.com In total, we hired 100 annotators, each annotating 20 examples (4 model outputs for each of the 5 domains). We selected annotators with at least 10 completed tasks and 100% approval rate, having English as their primary language. We paid the annotators £9 per hour, according to the platform’s recommendations. The median time for completing the annotations was 47 minutes. See Appendix C for the instructions for the annotators and the annotation interface and Appendix E for the data visualizations.
@@ 5 Results and Discussion
A summary of the token-level annotations is in Table 5 and 5, with detailed results per domain provided in Appendix F.
Table 3: The average numbers of errors per output (lower is better) based on GPT-4 ($\mathcal{E}_{\text{gpt}}$) and human annotators ($\mathcal{E}_{\text{hum}}$). We also include the average number of tokens per output in the rightmost column. The results of the best open LLM are emphasized.
Figure 5: The instructions given to the human annotators.
@@
$\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ Tok.
Llama2 2.46 1.50 0.90 1.21 0.20 0.23 0.13 0.10 3.70 3.05 83.8
Mistral 2.80 1.95 0.52 1.12 0.37 0.44 0.11 0.25 3.80 3.76 114.9
Zephyr 2.50 1.41 0.40 0.79 0.39 0.20 0.06 0.16 3.35 2.56 98.0
@@
$\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$
Llama2 0.78 0.52 0.46 0.57 0.15 0.17 0.09 0.08 0.92 0.85
Mistral 0.78 0.53 0.32 0.49 0.23 0.20 0.08 0.13 0.93 0.80
Zephyr 0.77 0.46 0.26 0.42 0.27 0.16 0.04 0.11 0.88 0.74
@@
Table 4: The ratio of outputs containing at least one error (lower is better) based on GPT-4 ($\mathcal{E}_{\text{gpt}}$) and human annotators ($\mathcal{E}_{\text{hum}}$). The results of the best open LLM are emphasized.
@@ 5.1 How Accurate Are the Model Outputs?
Depending on the model, between 74-85% of examples contain an error according to $\mathcal{E}_{\text{hum}}$, suggesting that open LLMs make semantic errors very often. According to $\mathcal{E}_{\text{gpt}}$, the number is as high as 88-93%.
The most common error type is INCORRECT. As shown in Table 5, all the open LLMs make more than two statements contradicting the data per output on average. The NOT_CHECKABLE errors are also relatively common: more than one per output on average according to $\mathcal{E}_{\text{hum}}$, and at least one being present in more than 26% of examples according to both metrics.
The results vary widely according to the domain (see Appendix F). For example, the outputs in wikidata contain much more NOT_CHECKABLE errors on average (1.54 per output according to $\mathcal{E}_{\text{hum}}$) than INCORRECT errors (0.12 per output according to $\mathcal{E}_{\text{hum}}$), suggesting that with simpler inputs, the models tend towards extrinsic hallucinations. The openweather domain seems to be the most complex with the longest outputs (~164 tokens), more than 8 errors in the output on average and >90% of outputs containing an error.
The differences between the open LLMs are not major. Out of the open LLMs, Zephyr has the best results across categories and metrics, followed by Llama2. However, the outputs of Mistral are longer on average, leaving more space for errors. GPT-3.5 (which we consider separately) does generally better according to both $\mathcal{E}_{\text{gpt}}$ and $\mathcal{E}_{\text{hum}}$, although still makes an error in 60-75% of examples (2 errors per example on average). In general, the results show that LLMs make too many semantic errors to be usable in practice for D2T generation in zero-shot setting.
@@ 5.2 Do Evaluation Methods Agree?
To quantify the agreement of our evaluation metrics, we computed the Pearson correlation coefficient between the error counts on the level of tokens, examples, and domains (see Appendix D for details). The correlation on the level of domains ($r_{\text{domain}}=0.92$) shows that $\mathcal{E}_{\text{gpt}}$ and $\mathcal{E}_{\text{hum}}$ correlate well for high-level ranking of the models. However, the correlation is weaker on the level of individual examples ($r_{\text{example}}=0.52$), and even weaker on the level of tokens ($r_{\text{token}}=0.25$).
In Table 5, we show the percentage of tokens marked by individual metrics. While both metrics mark around 21% of tokens as errorneous, they agree on the specific tokens in less than 6%, suggesting that the error spans may be subject to interpretation. We conclude while the details regarding error spans and categories may vary, the metrics generally agree on accuracy of model outputs for a given set of examples.
@@
Incorrect 0.135 0.099 0.040
Not checkable 0.039 0.076 0.016
Misleading 0.022 0.022 0.001
Other 0.008 0.018 0.001
All categories 0.204 0.214 0.059
Table 5: The ratio of tokens marked as errorneous by GPT-4 ($\mathcal{E}_{\text{gpt}}$), human annotators ($\mathcal{E}_{\text{hum}}$), and both metrics at the same time ($\mathcal{E}_{\text{gpt}}$ + $\mathcal{E}_{\text{hum}}$).
5.3 Recommendations and Directions
Forget fluency, solve accuracy.
The output of LLMs is mostly satisfactory regarding the style, format, and purpose of the text. However, the amount of semantic errors remains very high. Improving semantic accuracy of the models Li et al. (2022); Schmidtová (2023), along with new model-based evaluation metrics Liu et al. (2023); Xu et al. (2023), could thus help to bring improve LLM-based D2T generation systems where it is most needed.
Use efficient models.
The memory issues with long context, making few-shot experiments infeasible, can potentially be solved by using more efficient long-context models equipped with Flash Attention Dao et al. (2022) and fast inference libraries such as exllama [^6] https://github.com/turboderp/exllamav2 or llama.cpp. [^7] https://github.com/ggerganov/llama.cpp A potentially suitable open LLM in this respect is the long-context Llama2 (Xiong et al., 2023; model to be released).
@@ 5.3 Recommendations and Directions
Except for using an ad-hoc dataset of real-world data as we did in our work, ecological validity of D2T evaluation can be also ensured by continous evaluation with human users Zheng et al. (2023) and evaluating real-world impact of the systems Reiter (2023).
@@
Be careful about subtle bugs.
During our preliminary experiments, we uncovered subtle bugs in API calls such as incorrect instruction templates [^8] https://huggingface.co/docs/transformers/chat_templating or involuntary input truncation. With the apparent ease of API access and robustness of LLMs, such bugs could go unnoticed and artificially skew the model performance.
@@ 6.1 D2T Generation Tasks
Our phone specifications are closest to Wen et al. (2015, 2016) who introduced a dataset for generating descriptions of laptops and TVs. Their solution was based on recurrent neural networks, although templates remained a go-to approach for the task Wang et al. (2017). Recently, Shao et al. (2021) and Koto et al. (2022) also proposed specialized architectures based on pretrained language models for the data from big e-commerce platforms.
@@
The task of generating descriptions for a knowledge graph has been covered extensively in D2T generation (Gardent et al., 2017; Ferreira et al., 2020; Agarwal et al., 2021; Chen et al., 2020; Ribeiro et al., 2020, inter alia). Our task is to describe an entity provided a list of its properties, which is closely related to generating entity descriptions from Wikipedia infotables Lebret et al. (2016).
@@ 6.2 D2T Generation with LLMs
Recent works have focused on exploring capabilities of closed LLMs on existing D2T generation datasets. Axelsson and Skantze (2023) evaluated GPT-3.5 OpenAI (2023b) on WebNLG, along with Yuan and Färber (2023), who also tested the model on the AGENDA dataset Koncel-Kedziorski et al. (2019). Both works found out that regardless of potential data contamination, the LLMs rank behind state-of-the-art finetuned models on automatic metrics. Zhao et al. (2023) tested closed models on modified table-to-text generation datasets and found out that in terms of faithfulness, GPT-4 can outperform state-of-the-art models.
6.3 Beyond Reference-Based Metrics
Many works have recently investigated the potential of using LLMs for automatic reference-free evaluation of generated text, generally achieving high correlations with human judgement Zhao et al. (2023); Sottana et al. (2023); Kocmi and Federmann (2023a, b); Chiang and Lee (2023); Wang et al. (2023a); Fu et al. (2023). However, they also voice concerns about its non-reproducibility Kocmi and Federmann (2023a) and potential bias of these models Wang et al. (2023b).
Holtzman et al. (2023) suggest that the research on LLMs should move away from reporting benchmark scores, investigating model behaviors instead. In this vein, Upadhyay and Massie (2022) analysed the ability of models of producing different types of content in D2T generation. Regarding human evaluation, Thomson and Reiter (2020) proposed a protocol for reference-free token-level annotation of complex D2T generation output.
@@ 7 Conclusion
We provided an exploratory study into D2T generation with open LLMs. We proposed new directions for D2T generation, including using ad-hoc test sets, data in common formats, and reference-free evaluation. By a combination of GPT-4-based metric and human evaluation, we evaluated the performance of LLMs on five domains, providing token-level annotations of model outputs across five domains, and recommendations for future directions in D2T generation.
@@ Acknowledgements
This work was supported by the European Research Council (Grant agreement No. 101039303, NG-NLG) and Charles University project SVV 260 698. It used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No. LM2018101).
@@
In our work, we do not include comparison to other D2T generation approaches. The main reason is that our benchmark is reference-free, while large majority of prior approaches are based on models finetuned on reference outputs. However, we believe that our work still satisties our main goal of providing insights into behaviors of open LLM models on D2T generation.
We acknowledge that reference-free metrics currently have various shortcomings, including reliance on closed models or specific human annotation protocols, leading to limited replicability, and high price of execution. None of the approaches also produces flawless outcomes and have only moderate correlations with each other. We believe that these shortcomings will be adressed in the future with open model-based metrics.
@@
The approaches based on LLMs may produce factually incorrect information. Any text produced by the LLMs therefore need to be carefully examined, and no decisions should be based on the generated text alone.

### v3

@@ 4.2 GPT-4-based Evaluation
For automatic evaluation, we leverage the fact that LLM-based metrics can be customized for a particular task without the need for training data. In our experiments, we use a metric based on GPT-4 (gpt-4-1106-preview, OpenAI, 2023a), which was shown to be superior to other LLMs in following fine-grained instructions, reaching high correlations with human judgment on evaluating generated texts Zhao et al. (2023); Sottana et al. (2023); Kocmi and Federmann (2023a, b). [^10] We confirmed that GPT-3.5 and Llama 3 have lower correlations with human judgments also in our scenario, see Appendix D.
We instantiate $\mathcal{E}_{\text{gpt}}$ with a prompt and a system message describing the task. We instruct the model to produce a JSON output with sequentially ordered errors using the following format:
@@
Note that we require that the model first generates the free-form text reason for the error. Generating the reason comes at almost no extra cost and our cursory observations suggest that requiring it leads to more precise outputs. [^11] We did not ask the crowdworkers for free-form reasoning about the errors since that would make the annotation notably more complex.
We align the model outputs with the original text by string matching on TEXT_SPAN, moving the current position forward after each match. We ensure that the model response is a valid JSON using OpenAI’s response_format parameter. See Appendix C for more details about the metric, including the prompt and the system message.
@@ 5 Results and Discussion
A summary of the word-level annotations is in section 5 and 5, with detailed results per domain provided in Appendix F.
@@
$\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ Tok.
Llama 2 1.57 2.79 1.25 0.91 0.25 0.12 0.10 0.09 3.18 3.90 83.8
Mistral 2.03 3.23 1.12 0.54 0.44 0.26 0.25 0.10 3.85 4.12 114.9
Zephyr 1.44 2.84 0.77 0.40 0.20 0.29 0.16 0.05 2.58 3.58 98.0
@@
Table 3: The average numbers of errors per output (lower is better) based on human annotators ($\mathcal{E}_{\text{hum}}$) and GPT-4 ($\mathcal{E}_{\text{gpt}}$). We also include the average number of tokens per output in the rightmost column. The results of the best open LLM are emphasized.
@@
$\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$
Llama 2 53.2% 80.0% 57.4% 44.8% 17.4% 8.8% 7.6% 7.6% 85.6% 94.0%
Mistral 53.6% 80.2% 49.6% 31.8% 20.6% 17.0% 13.6% 8.4% 81.2% 93.0%
Zephyr 46.8% 78.0% 42.2% 25.0% 16.2% 20.6% 11.6% 4.2% 75.6% 89.4%
@@
Table 4: The percentage of outputs containing at least one error (lower is better) based on human annotators ($\mathcal{E}_{\text{hum}}$) and GPT-4 ($\mathcal{E}_{\text{gpt}}$). The results of the best open LLM are emphasized.
@@ 5.1 How Accurate Are the Model Outputs?
Depending on the model, between 76-86% of examples contain an error according to $\mathcal{E}_{\text{hum}}$, suggesting that open LLMs make semantic errors very often. According to $\mathcal{E}_{\text{gpt}}$, the number is as high as 89-94%.
The most common error type is INCORRECTI. As shown in section 5, all the open LLMs make more than two statements contradicting the data per output on average. The NOT_CHECKABLENC errors are also relatively common: more than one per output on average according to $\mathcal{E}_{\text{hum}}$, and at least one being present in more than 25% of examples according to both metrics.
The results vary widely according to the domain (see Appendix F). For example, the outputs in wikidata contain much more NOT_CHECKABLENC errors on average (1.01 per output according to $\mathcal{E}_{\text{hum}}$) than INCORRECTI errors (0.11 per output according to $\mathcal{E}_{\text{hum}}$), suggesting that with simpler inputs, the models tend to introduce extra information. The openweather domain seems to be the most complex with the longest outputs (~164 tokens), more than eight errors in the output on average, and >90% of outputs containing an error.
The differences between the open LLMs are not major. Out of the open LLMs, Zephyr has the best results across categories and metrics, followed by Llama 2. However, the outputs of Mistral are longer on average, leaving more space for errors. GPT-3.5 (which we consider separately) does generally better according to both $\mathcal{E}_{\text{gpt}}$ and $\mathcal{E}_{\text{hum}}$, although it still makes an error in 60-75% of examples (2 errors per example on average). In general, the results show that LLMs make too many semantic errors to be usable in practice for D2T generation in a zero-shot setting.
@@ 5.2 Do Evaluation Methods Agree?
To quantify the agreement of $\mathcal{E}_{\text{hum}}$ and $\mathcal{E}_{\text{gpt}}$, we computed the Pearson correlation coefficient between the error counts on the level of words, examples, and domains as follows (note that each error category was considered separately):
  • For $r_{\text{domain}}$, we used the average error counts per domain (see section 5).
  • For $r_{\text{example}}$, we used the count of errors per example.
  • For $r_{\text{word}}$, we used the binary indicators marking an error per word.
The correlation on the level of words is weak ($r_{\text{word}}=0.26$) but gets better on the example-level ($r_{\text{example}}=0.52$) and even better on the domain-level ($r_{\text{domain}}=0.93$). In Table 5, we show the percentage of words marked as errors by individual metrics. The metrics agree on the specific words in less than 6%, although they both mark around 21% of words as erroneous.
We also measure inter-annotator agreement between human annotators. For that, we obtained annotations from two annotators for 100 model outputs. The results are similar: the annotators agree weakly on the word level ($r_{\text{word}}=0.36$), stronger on the example level ($r_{\text{example}}=0.53$), and even stronger on the domain level ($r_{\text{domain}}=0.85$). We conclude that while the details regarding error spans and categories may vary, the annotators as well as GPT-4 generally agree on the accuracy of model outputs for a given set of examples. In the future, the agreement could be improved by measuring errors on the phrase level Vamvas and Sennrich (2022).
@@
Incorrect 10.1% 14.2% 4.1%
Not checkable 7.8% 4.3% 2.0%
Misleading 2.2% 1.5% 0.1%
Other 1.8% 0.7% 0.1%
Total 21.9% 20.7% 6.3%
Table 5: The percentage of words marked as erroneous by human annotators ($\mathcal{E}_{\text{hum}}$), GPT-4 ($\mathcal{E}_{\text{gpt}}$), and both approaches at the same time ($\mathcal{E}_{\text{hum}}$ + $\mathcal{E}_{\text{gpt}}$).
5.3 Recommendations for Future Work
Focus on semantic accuracy.
The output of LLMs is satisfactory regarding the style, format, and purpose of the text. However, the amount of semantic errors remains very high. Improving the semantic accuracy of the models Li et al. (2022), along with new model-based evaluation metrics Liu et al. (2023); Xu et al. (2023), could thus help to bring improve LLM-based D2T generation systems where it is most needed.
Use efficient and long-context models.
The memory issues with long context, making few-shot experiments infeasible, can potentially be solved by using more efficient models equipped with Flash Attention Dao et al. (2022) and fast inference libraries such as llama.cpp [^12] https://github.com/ggerganov/llama.cpp. The recent emergence of capable long-context models Bai et al. (2023); Munkhdalai et al. (2024) also seems to play in favor of LLM-based D2T generation with long inputs.
Be careful about subtle bugs.
During our preliminary experiments, we fixed several subtle bugs in our API calls such as incorrect instruction templates [^13] https://huggingface.co/docs/transformers/chat_templating or involuntary input truncation. Therefore, we recommend careful checks of API calls, as with the apparent ease of API access and robustness of LLMs, such bugs could go unnoticed and artificially skew the model performance.
@@ 5.3 Recommendations for Future Work
Except for using an ad-hoc dataset of real-world data as we did in our work, the ecological validity of D2T evaluation can also be ensured by continuous evaluation with human users Zheng et al. (2023) and evaluating the real-world impact of the systems Reiter (2023).
@@ 6 Related Work
6.1 Evaluation of Generated Texts
Evaluation of generated texts is a complex task lacking a generally accepted solution Celikyilmaz et al. (2020). While researchers are acknowledging the importance of combining multiple evaluation metrics Gehrmann et al. (2021, 2023), most evaluation is still based on comparing the model outputs to human-written references, which tend to be noisy and expensive to collect Dušek et al. (2019, 2020).
Many works recently investigated the potential of using LLMs for automatic reference-free evaluation of generated texts, generally achieving high correlations with human judgment Zhao et al. (2023); Sottana et al. (2023); Kocmi and Federmann (2023a, b); Chiang and Lee (2023); Wang et al. (2023a); Fu et al. (2023). However, they also voice concerns about its non-reproducibility Kocmi and Federmann (2023a) and potential bias of these models Wang et al. (2023b).
Human evaluation is an essential component of natural language generation experiments van der Lee et al. (2019, 2021). The closest human evaluation protocol to our scenario is the reference-free word-level annotation of complex D2T generation output proposed in Thomson and Reiter (2020) and Thomson et al. (2023).
@@ 6.2 D2T Generation Tasks
Our phone specifications are closest to Wen et al. (2015, 2016), who introduced a dataset for generating descriptions of laptops and TVs. Their solution was based on recurrent neural networks, although templates remained a go-to approach for the task Wang et al. (2017). Recently, Shao et al. (2021) and Koto et al. (2022) also proposed specialized architectures based on pretrained language models for the data from big e-commerce platforms.
@@
The task of generating descriptions for a knowledge graph has been covered extensively in D2T generation (Gardent et al., 2017; Castro Ferreira et al., 2020; Agarwal et al., 2021; Chen et al., 2020; Ribeiro et al., 2020, inter alia). Our task is to describe an entity provided a list of its properties, which is closely related to generating entity descriptions from Wikipedia infotables Lebret et al. (2016).
@@ 6.3 D2T Generation with LLMs
Recent works have focused on exploring the capabilities of closed LLMs on existing D2T generation datasets. Axelsson and Skantze (2023) evaluated GPT-3.5 OpenAI (2023b) on WebNLG, along with Yuan and Färber (2023), who also tested the model on the AGENDA dataset Koncel-Kedziorski et al. (2019). Both works found that regardless of potential data contamination, the LLMs rank behind state-of-the-art finetuned models on automatic metrics. Zhao et al. (2023) tested closed models on modified table-to-text generation datasets and found out that in terms of faithfulness, GPT-4 can outperform state-of-the-art models.
@@ 7 Conclusion
We provided an exploratory study into D2T generation with open LLMs. We proposed new directions for D2T generation, including using ad-hoc test sets, data in common formats, and reference-free evaluation. By a combination of GPT-4-based metric and human evaluation, we evaluated the performance of LLMs on five domains, providing word-level annotations of model outputs across five domains and recommendations for future directions in D2T generation.
@@ Acknowledgements
This work was funded by the European Union (ERC, NG-NLG, 101039303) and Charles University project SVV 260 698. It used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No. LM2018101).
@@
In our work, we do not include a comparison to other D2T generation approaches. The main reason is that our benchmark is reference-free, while a large majority of prior approaches are based on models finetuned on reference outputs. However, we believe that our work still satisfies our main goal of providing insights into behaviors of open LLM models on D2T generation.
We acknowledge that reference-free metrics currently have various shortcomings, including reliance on closed models and specific human annotation protocols, leading to limited replicability and a high price of execution. The approaches occasionally produce incorrect outcomes themselves and they have only moderate correlations with each other. We believe that these shortcomings will be addressed in the future with open model-based metrics.
@@
The approaches based on LLMs may produce factually and semantically incorrect information. Any text produced by the LLMs therefore needs to be carefully examined, and no decisions should be based on the generated text alone. Discovering the causes of LLMs’ hallucinations is out of scope of this paper, but is currently a major topic under investigation Ji et al. (2023).
Ethical Considerations
The human evaluation study was approved by the internal ethics committee of our institution. The annotators were hired over Prolific and paid the platform-recommended wage of 9 GBP/hour. The annotators were preselected based on their primary language (English) and their country of residence (US, UK, Ireland, Australia, New Zealand). All annotators were shown detailed instructions and explanation of the data types, data sources, and the purpose of the research (see Appendix B for details). The domains in Quintd were selected so that they do not contain any sensitive or potentially offensive content. We do not collect any demographic data about the participants.